OpenStudy (anonymous):

I am looking for the number of neutrons of a second isotope. The specific charge = total charge of nucleus/mass of nucleus = 3.7x10^7 total charge of nucleus = 92 x 1.6x10^-19 mass of nucleus = A x 1.67x10^-27 where, A = number of neutrons + number of protons You need to rearrange and find A, then subtract the number of protons, the answer should be about 146 neutrons. Can anyone help?

OpenStudy (unklerhaukus):

the charge of the nucleus is entirely due to the number of protons The charge in coulombs is \(92 \times 1.6\times10^{-19} C\) \(1.6\times10^{-19} C\) is the magnitude of the charge of an electron or a proton divide the total charge of the nucleus by the charge of the proton to find the number of protons in the nucleus \[Z=\frac{92 \times 1.6\times10^{-19} C}{1.6\times10^{-19} C} \]

OpenStudy (unklerhaukus):

\[A=Z+N\]

OpenStudy (anonymous):

That equals 92?

OpenStudy (unklerhaukus):

yes 92 protons, so that can only correspond to one element right, can you find the element on a periodic table,?

OpenStudy (anonymous):

Yes that is uranium?

OpenStudy (unklerhaukus):

that is right \(^A_{92}U\). the mass of the nucleus in kilograms \(m=A \times 1.67\times10^{-27} kg\) the atomic mass unit \[u=1.67\times 10^{-27} kg\]

OpenStudy (anonymous):

Ok whats next?

OpenStudy (unklerhaukus):

you should get \(m=Au\) simply dived by \(u\) to get \(A\)

OpenStudy (unklerhaukus):

and set the two equations for \(A\) equal to one another , all that is left is to solve for the number of neutrons \(N\)

OpenStudy (anonymous):

Yes i know. The number of neutrons is 146. I already have the answer. 146=92=238! BUT to get 146 you have to come to the calculation of 238 how do i solve that?

OpenStudy (unklerhaukus):

total charge of nucleus/mass of nucleus \(=Ze/m\)

OpenStudy (anonymous):

Ok. That is 92 *e/m?

OpenStudy (unklerhaukus):

\[\frac{Ze}{m}=\frac{Ze}{Au}=\frac ZA\frac eu=3.7\times10^7 (C/kg)\]

OpenStudy (anonymous):

Yes that,s correct 3.7*10^7 (C/kg). What next?

OpenStudy (unklerhaukus):

solve for \(a\) remember \(a\) should be 238

OpenStudy (anonymous):

Yes i know. How do we solve A?

OpenStudy (unklerhaukus):

i ment capital \(A\) not \(a\)

OpenStudy (anonymous):

Ok A

OpenStudy (unklerhaukus):

\[A=\frac {Ze}{3.7\times10^{7}u} \]

OpenStudy (anonymous):

Yes that = 3.97*10-25 is that correct?

OpenStudy (unklerhaukus):

im getting confused

OpenStudy (anonymous):

Ok. 92*1.6*10^7/3.7*10^7 =3.97*10^-25. Or am i doing something wrong here?

OpenStudy (unklerhaukus):

no that is good , my mistake \[A=\frac{Ze}{qu}=\frac{92\times1.6\times10^{-19}}{3.7\times10^7u}=\frac{3.97\times10^{-25}}{u} \rightarrow \frac{3.97\times10^{-25}}{1.67×10^{−27}kg}=\dots \]

OpenStudy (unklerhaukus):

i kinda messed up the units a bit though

OpenStudy (anonymous):

=237.72 or 238. Yes!Yes!Yes thank god someone gets it at last. Well done

OpenStudy (unklerhaukus):

\[A=\frac{Z\cdot e[C]}{q[C/kg]\cdot u[kg]}=\frac{92~\cdot~1.6\times10^{-19}[C]}{3.7\times10^{7}~\cdot~ 1.67\times10^{-27}[1/C]}=238\]

OpenStudy (anonymous):

Well done. You deserve more than one medal for persevering with it!

OpenStudy (unklerhaukus):

And we have our answer \[^{238}_{92}U\] unfortunatly not as useful as \(^{235}_{92}U\) but at least it will last longer

OpenStudy (anonymous):

Yes. Till the next time